# Decoding Cortical Microcircuits: A Generative Model for Latent Space Exploration and Controlled Synthesis

### v1

@@ Abstract
A central idea in understanding brains and building artificial intelligence is that structure determines function. Yet, how the brain’s complex structure arises from a limited set of genetic instructions remains a key question. The ultra high-dimensional detail of neural connections vastly exceeds the information storage capacity of genes, suggesting a compact, low-dimensional blueprint must guide brain development. Our motivation is to uncover this blueprint. We introduce a generative model, to learn this underlying representation from detailed connectivity maps of mouse cortical microcircuits. Our model successfully captures the essential structural information of these circuits in a compressed latent space. We found that specific, interpretable directions within this space directly relate to understandable network properties. Building on this, we demonstrate a novel method to controllably generate new, synthetic microcircuits with desired structural features by navigating this latent space. This work offers a new way to investigate the design principles of neural circuits and explore how structure gives rise to function, potentially informing the development of more advanced artificial neural networks.
1 Introduction
The relationship between structure and function is a core idea in both neuroscience and the development of artificial intelligence Gradientsofstructure-functiontetheringacrossneocortex ; NAS1 ; NAS2 . The brain, with its amazing abilities, inspires us to build better AI systems neuronAI1 ; neuronAI2 ; hassabis2017 ; lecun2015 . However, the brain’s structure is incredibly complex complex_connectome1 ; complex_connectome2 . Understanding how all its neurons are connected is a huge task. These maps of connections are called connectomes connectome_def1 ; connectome_def2_original ; connectome_def3 . Studying connectomes is important because they hold clues about how the brain processes information and learns HCP ; USbrain1 .
@@ 1 Introduction
A key puzzle is how the brain’s intricate wiring instructions are stored. An animal’s genes guide the development of its nervous system Genetic_influences_on_hub_connectivity_of_the_human_connectome ; Genetic_architecture_of_the_structural_connectome . However, the amount of information genes can hold is much smaller than what would be needed to explicitly list every single connection in a fully formed brain info_capacity1 ; info_capacity2 . This observation, known as the "genetic bottleneck" pnas_genetic_btnk , suggests that there must be a simpler, more compact set of rules or a low-dimensional representation that guides how brain networks grow and organize themselves.
@@
This paper aims to uncover such a low-dimensional code for brain circuitry. We focus on cortical microcircuits, which are small, repeating patterns of cortical connections that can be thought of as fundamental building blocks of the brain. To do this, we use data from the MICrONS program microns , which has produced a large and detailed connectome dataset from the visual cortex of a mouse. This dataset provides an unprecedented opportunity to study the organization of these microcircuits.
@@
We name our model NeuroStructGen, a Neural Structure Generator. To our knowledge, this combined approach of learning a compact representation of biological connectomes and then using its understandable features for precise, controllable generation of new circuits has not been done before. This work offers a new way to explore how changes in basic organizational rules might affect brain circuit structure. Furthermore, understanding the brain’s structure could eventually help in designing more efficient and capable artificial neural networks.
2 Problem Definition
The core motivation of this work stems from the “genetic bottleneck” hypothesis: the vast complexity of the brain’s connectome cannot be explicitly encoded by the genome, suggesting the existence of a compact, low-dimensional generative code. Our overarching goal is to develop a computational framework to learn, interpret, and control such generative representations of brain microcircuit topology. We define two primary objectives:
2.1 Learning Compact Generative Representations of Microcircuit Topology
Given a dataset of $N$ biological microcircuit graphs, denoted as $\mathbf{G}_{\text{data}}=\{\mathcal{G}^{(1)},\mathcal{G}^{(2)},\dots,\mathcal{G}^{(N)}\}$, the first objective is to learn a low-dimensional latent representation $\mathbf{z}^{(i)}\in\mathbb{R}^{d_{z}}$ for each microcircuit $\mathcal{G}^{(i)}$. Each microcircuit is represented as a graph $\mathcal{G}^{(i)}=(\mathbf{X}^{(i)},\mathbf{A}^{(i)})$, where $\mathbf{X}^{(i)}\in\mathbb{R}^{n_{i}\times d_{v}}$ is a matrix of node features for its $n_{i}$ neurons (with $d_{v}$ feature dimensionality), and $\mathbf{A}^{(i)}\in\{0,1\}^{n_{i}\times n_{i}}$ is its adjacency matrix indicating synaptic connections. In this work, since the number of nodes varies across the graph data, we pad each adjacency matrix to 100×100. For consistent processing, a canonical node ordering $\pi$ is assumed for inputs to certain model components, thus we may refer to an ordered graph as $\mathcal{G}_{\pi}^{(i)}$.
@@ 2.1 Learning Compact Generative Representations of Microcircuit Topology
2.2 Controllable Generation of Microcircuits with Target Properties
Building upon the learned latent space $\mathcal{Z}$ (the space of all $\mathbf{z}$) and the generative model $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$ from Section 2.1, our second objective is to enable the controlled generation of novel microcircuits.
Specifically, given a target structural, dynamical, or functional property $\mathcal{T}$ (e.g., a specific mean degree, clustering coefficient, or level of assortativity), and a desired value $t_{\text{target}}$ for this property, the goal is to synthesize new connectome graphs $\mathcal{G}_{\pi,\text{new}}$. These generated graphs should:
@@ 2.2 Controllable Generation of Microcircuits with Target Properties
    Optimally satisfy the specified constraint, meaning the property $\mathcal{T}$ for $\mathcal{G}_{\pi,\text{new}}$ should be close to $t_{\text{target}}$.
@@
Formally, this requires effectively sampling from, or being guided by, a conditional probability distribution $p(\mathbf{z}|\mathcal{T},t_{\text{target}})$ in the latent space. Latent vectors $\mathbf{z}_{\text{new}}$ drawn from this distribution are then decoded using $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z}_{\text{new}})$ to produce the desired microcircuits.
3 Related Work
@@
Analyzing the structural organization of brain connectomes is fundamental to understanding the intricate relationship between brain structure and function ref1 ; ref2 . A number of methods have been used to address this link, including statistical models statistical_models , communication models communication_model1 , and biophysical models biophysical_models1 ; biophysical_models2 . By modeling brain networks as graphs, researchers employ graph-theoretic approaches to reveal key topological properties that underlie efficient communication and cognitive processes ref3 ; ref4 . Understanding how these structural features relate to functional dynamics is a central goal in connectomics.
@@
To investigate the principles governing connectome formation and organization, generative models have been developed. Early approaches typically relied on a small set of predefined wiring rules or biological principles, such as cost-efficiency or growth mechanisms, to replicate observed network features in silico Generativemodelsofthehumanconnectome ; Modellingthedevelopmentofcorticalsystemsnetworks ; ASimpleGenerativeModel . While successful in capturing certain global properties, these rule-based methods often lack flexibility and depend on manually specified or constrained generative factors, limiting their ability to explore the full complexity of biological variability.
@@
More recently, deep generative models, such as Variational Autoencoders (VAEs), have emerged as powerful tools for analyzing and synthesizing complex data like brain networks ref5 ; ref6 . These models are particularly well-suited for learning a low-dimensional latent representation of the network data, effectively performing dimensionality reduction. Furthermore, by manipulating or sampling from this learned latent space, these models enable the controlled synthesis of novel, biologically plausible network configurations, offering new avenues for exploring connectome variability and its functional implications ref7 ; ref8 .
4 Method
4.1 MICrONS Dataset and Preprocess
@@
We adopt the IARPA MICrONS dataset microns , which encompasses a 1.4 × 0.87 × 0.84 mm volume of cortex from a P87 mouse microns , containing neurons within the area and their interconnections. Given the cortex’s division into six distinct layers, each associated with a specific level of information processing layers1 ; layers2 ; layer3 , our study concentrates on cortical microcircuits organized as vertically oriented functional columns that traverse all laminar layers circuit1 ; circuit2 . To model these microcircuits, we extracted cylindrical subvolumes oriented perpendicularly to the cortical layers. Each cylindrical unit comprises 80–100 neurons, preserving intra-column connectivity while excluding connections extending beyond the columnar boundary.
Each columnar circuit is treated as a binary directed graph. To maintain the relative positional relationship between neurons, we define an ordering rule $\pi$ based on their y-coordinates, representing their positions within the cortical hierarchy. This consistent sorting rule also facilitates the comparison between the original data and generated results during the training and evaluation of the generative model. The relative positions of neurons are encoded by one-hot vectors indicating their indices according to the defined order $\pi$. Neuron type and synapse weight information are temporarily disregarded, as this work focuses solely on the topological structure.
4.2 Connectome Graph Variational Autoencoder
@@
As the goal is to find latent representations similar to “genetic bottleneck”, we need a generative model with “information bottleneck”. Thus, adopting a variational autoencoder (VAE) VAE on the connectome data can be formulated as follows: Assume a set of training connectome graphs $\mathbf{G}=\{\mathcal{G}_{\pi}^{(i)}\}$ is generated from the distribution of a set of unobserved latent representation $\textbf{z}=(z_{1},...,z_{m})$, where $z^{(i)}\sim p_{\theta^{*}}(z)$ and training data is sampled from true conditional $p_{\theta^{*}}(\mathcal{G}_{\pi}|z^{(i)})$. We denote the data generation process as $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$. Following the standard VAE setting VAE , we approximate the intractable posterior by $q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi})\approx p_{\theta}(\mathbf{z}|\mathcal{G}_{\pi})$ and minimize the evidence lower bound on the marginal likelihood of graph $\mathcal{G}^{(i)}$:
@@
$\log p_{\theta}(\mathcal{G}_{\pi}^{(i)})\geq\mathcal{L}(\phi,\theta;\mathcal{G}_{\pi}^{(i)})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi}^{(i)})}\left[\log p_{\theta}(\mathcal{G}_{\pi}^{(i)}|\mathbf{z})\right]-\beta\mathrm{KL}\left[q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi}^{(i)})||p_{\theta}(\mathbf{z})\right],$ (1)
@@
where $\beta$ is a hyperparameter to balance the reconstruction loss and KL-divergence loss during the training process betaVAE .
@@
The proposed VAE model for connectome graphs comprises four main components. First, a node feature encoder employs a multi-layer, multi-head Graph Attention Network (GAT) GAT ; attentionisallyouneed to compute an embedding vector for each node. Second, a graph global encoder, which is a transformer encoder augmented with a special token BERT ; pigvae ; attentionisallyouneed , extracts an embedding for the entire sequence of node embeddings. This extracted sequence embedding serves as the graph-level embedding, upon which a 32-dimensional mean and variance are computed. The reparameterization of a standard VAE is applied to this 32D latent embedding representing the whole graph. Third, a node feature decoder, which is a transformer decoder, takes the graph-level embedding as input to reconstruct node features necessary for subsequent edge prediction. Finally, the edge predictor utilizes these decoded node features to predict the edges of the graph.
@@
The overall architecture of the VAE model is illustrated in Figure 1. The detailed encoder structure is depicted in Appendix Figure 8, and the detailed decoder structure is depicted in Appendix Figure 9.
@@
Node Feature Encoder
The node feature encoder, utilizing a three-layer multi-head GAT network, transforms 100-dimensional one-hot node representations into 32-dimensional node embeddings. See Appendix 7.2 for GAT architecture details.
Graph Global Encoder
Treating the y-ordered nodes as a sequence (analogous to words in a sentence), the graph global encoder transforms node embeddings into a fixed-size latent representation. Following BERT ; pigvae , a dummy node $v_{0}$ is prepended to the sequence to serve as a global embedding. The encoder applies rotational positional encoding (RoPE) rope and several transformer encoder layers to this augmented sequence. The final embedding of $v_{0}$ is taken as the global graph representation. An MLP is then applied to compute the 32-dimensional mean and variance for sampling the 32-dimensional latent vector $\mathbf{z}$, similar to standard VAEs.
Node Feature Decoder
The Node Feature Decoder reconstructs individual node embeddings from the global graph embedding using several transformer decoder layers. The input is the global graph embedding, augmented with rotational positional encoding (RoPE) rope . The global graph embedding also serves as memory for the decoder’s cross-attention, leveraging this global context during node feature reconstruction.
Edge Predictor
4.3 Controllable Connectome Generation by Sampling from the Latent Space
@@
When $\lambda\to\infty$, the energy model degenerate into a strictly constrained version (by $\mathcal{T}$). In this case we have:
@@
[…]
5 Experiments
5.1 Evaluation Metrics
We selected several graph theory metrics (descriptors) commonly employed in the analysis of connectomes brain_graph_metrics , including mean degree, efficiency, transitivity, clustering coefficient, modularity, and assortativity. The detailed definitions of these metrics are provided in the Appendix 7.3.
5.2 Microcircuit Reconstruction and Generation Results
Examples of original and reconstructed graphs obtained through the VAE are provided in the Appendix 7.4. To validate the model’s generative capabilities, we evaluated several graph metrics on the generated samples. We employed the maximum mean discrepancy (MMD) MMD to compare the distributions of these graph statistics between an equal number of generated and test graphs. Following the methodology in you2018b_mmd , we specifically measured the distributions of degree, clustering coefficient and spectrum. For MMD computation, we utilized both the Gaussian Earth Mover’s Distance (EMD) kernel and the total variation (TV) distance liao2019 . Given the absence of existing generative models specifically designed for connectome graphs, we selected three alternative models commonly used for molecule generation or general graph generation GDSS ; DisCo ; EDGE as baselines. The following Table 1 reports the MMD values for different graph metrics evaluated across these models and our proposed approach. Figure 2 compares graphs generated by our proposed model and other models with the original data (column Ori.).
@@ 5.2 Microcircuit Reconstruction and Generation Results
GDSS 1.138 0.493 1.381 0.959 0.325 0.515
DisCo 1.047 0.304 1.315 0.550 0.094 0.334
EDGE 0.660 0.041 0.343 1.016 0.972 0.111
NeuroStructGen (Ours) $\boldsymbol{0.164}$ $\boldsymbol{0.028}$ $\boldsymbol{0.154}$ $\boldsymbol{0.021}$ $\boldsymbol{0.047}$ $\boldsymbol{0.007}$
Table 1: Comparison with Baseline Models
@@
Figure 2: Generated result comparison.
5.3 Latent Representation Analysis
5.3.1 SHAP Analysis

### v2

@@ Abstract
A central idea in understanding brains and building brain neural networks is that structure determines function. However, the brain’s connectome is a massively high-dimensional graph, making the direct investigation of its structure-function relationships computationally intractable. Therefore, identifying a compact, low-dimensional representation that captures the connectome’s essential structural organization is crucial for elucidating these relationships. The existence of such a representation is biologically plausible: the ”genomic bottleneck” theory provides a strong basis for such a compressed developmental blueprint. We introduce a generative model to learn this underlying representation from detailed connectivity maps of mouse cortical microcircuits. Our model successfully captures the essential structural information of these circuits within a compressed latent space. We then associate specific network structures, as encoded in this latent space, with computational functions using reservoir computing tasks. Building on this, our methodology allows for the controllable generation of novel, synthetic microcircuits with desired structural features by navigating the learned latent space. This research paradigm establishes a computational testbed to systematically investigate the brain’s inherent structure-function relationships. The ability to generate diverse, bio-plausible circuits could inform the development of more advanced artificial neural networks.
Introduction
The relationship between structure and function is a core idea in both neuroscience and the development of artificial intelligence (Vázquez-Rodríguez et al. 2019; Clark 2023; White et al. 2023). The brain, with its extraordinary capabilities, inspires us to build better AI systems (Zador et al. 2023; Osegi 2023; Hassabis et al. 2017; Lecun, Bengio, and Hinton 2015). The maps of brain connections are called connectomes (Sporns 2013; Sporns, Tononi, and Kötter 2005; Sporns 2011), whose structure is incredibly complex (Craddock, Tungaraza, and Milham 2015; Collin and Whitfield-Gabrieli 2023). Studying connectomes is important because they hold clues about how the brain processes information and learns (Elam et al. 2021; Bargmann and Newsome 2014). However, given that the brain’s connectome is a massively high-dimensional graph, it is impractical to directly investigate which structures within the connectome itself yield specific functional performance. Therefore, the objective of this paper is to find a low-dimensional representation that captures the key features of the connectome. This new paradigm leverages this low-dimensional representation as a novel tool to investigate structure-function relationships.
@@ Introduction
The existence of such a low-dimensional representation is also supported by the ’genomic bottleneck’ theory. An animal’s genes guide the development of its nervous system (Arnatkevičiute et al. 2020; Wainberg et al. 2024). However, the amount of information genes can hold is much smaller than what would be needed to explicitly list every single connection in a fully formed brain (Suganuma et al. 2020; Nigam et al. 2024). This observation, known as the ’genomic bottleneck’ (Shuvaev et al. 2024), suggests (without negating the combined influence of genetics and environment) that there must be a simpler, more compact set of rules or a low-dimensional representation that guides how brain networks grow and organize themselves, for which the finite information capacity of the genome provides primary evidence.
@@
In this work, we focus on cortical microcircuits, which are small, repeating patterns of cortical connections that can be thought of as fundamental building blocks and basic computational units of the brain (Douglas and Martin 2004a; Miller 2016). We primarily implement this new paradigm for investigating structure-function relationships using data from the MICrONS program (Consortium et al. 2023), which has produced a large and detailed connectome dataset from the visual cortex of a mouse. This dataset provides an unprecedented opportunity to study the organization of these microcircuits. As a supplementary study, we also conduct similar investigations on the FlyWire connectome dataset from the fruit fly (See Appendix L).
@@
    We propose a new paradigm for investigating the structure-function relationships of the brain connectome, which we implement on connectome data from both mice and fruit flies. While the core paradigm remains consistent, the specific datasets and model architectures differ between species. Due to space constraints, this paper primarily details the model for the mouse connectome, with results from the fruit fly experiments presented in the Appendix L.
  2. 2.
@@
  5. 5.
    We investigated the influence of connectivity patterns on network function across various tasks. Utilizing controllably generated networks, we constructed reservoir networks and demonstrated that networks generated by the VAE to emulate brain-like connectivity patterns exhibited enhanced task performance compared to randomly connected networks of the same density. Furthermore, we found that alterations in specific structural features had an approximately monotonic effect on the performance of certain tasks.
To our knowledge, the application of VAE for controllable generation of high-resolution structural microcircuits and for the understanding of structural-function relationship has not been done before. Furthermore, understanding the brain’s structure could eventually help in designing more efficient and capable artificial neural networks.
Problem Definition
Our overarching goal is to develop a computational framework that learns a low-dimensional generative representation of brain microcircuit topology. This representation is intended to serve as a new paradigm for investigating structure-function relationships, providing a basis for controllably generating novel microcircuits with desirable properties. We define two primary objectives:
Learning Compact Generative Representations of Microcircuit Topology
Given a dataset of $N$ biological microcircuit graphs, denoted as $\mathbf{G}_{\text{data}}=\{\mathcal{G}^{(1)},\mathcal{G}^{(2)},\dots,\mathcal{G}^{(N)}\}$, the first objective is to learn a low-dimensional latent representation $\mathbf{z}^{(i)}\in\mathbb{R}^{d_{z}}$ for each microcircuit $\mathcal{G}^{(i)}$. Each microcircuit is represented as a graph $\mathcal{G}^{(i)}=(\mathbf{X}^{(i)},\mathbf{A}^{(i)})$, where $\mathbf{X}^{(i)}\in\mathbb{R}^{n_{i}\times d_{v}}$ is a matrix of node features for its $n_{i}$ neurons (with $d_{v}$ feature dimensionality), and $\mathbf{A}^{(i)}\in\{0,1\}^{n_{i}\times n_{i}}$ is its adjacency matrix indicating synaptic connections. In this work, since the number of nodes varies across the graph data, we pad each adjacency matrix to a fixed size of 100×100. For consistent processing, a canonical node ordering $\pi$ is assumed for inputs to certain model components, thus we may refer to an ordered graph as $\mathcal{G}_{\pi}^{(i)}$.
@@
Controllable Generation of Microcircuits with Target Properties
Building upon the learned latent space $\mathcal{Z}$ (the space of all $\mathbf{z}$) and the generative model $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$ from Section Learning Compact Generative Representations of Microcircuit Topology, our second objective is to enable the controlled generation of novel microcircuits.
Specifically, given a target structural, dynamical, or functional property $P$ (e.g., a specific mean degree, clustering coefficient, or level of assortativity), and a desired value $t_{\text{target}}$ for this property, the goal is to synthesize new connectome graphs $\mathcal{G}_{\pi,\text{new}}$. These generated graphs should:
@@
    Optimally satisfy the specified constraint, meaning the property $P$ for $\mathcal{G}_{\pi,\text{new}}$ should be close to $t_{\text{target}}$. We denote this constraint as $\mathcal{T}$.
@@
Formally, this requires effectively sampling from, or being guided by, a conditional probability distribution $p(\mathbf{z}|\mathcal{T})$ in the latent space. Latent vectors $\mathbf{z}_{\text{new}}$ drawn from this distribution are then decoded using $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z}_{\text{new}})$ to produce the desired microcircuits.
Related Work
@@
Analyzing the structural organization of brain connectomes is fundamental to understanding the intricate relationship between brain structure and function (Xia, Wang, and He 2013; Chen et al. 2018). A number of methods have been used to address this link, including statistical models (Mišić et al. 2016), communication models (Goñi et al. 2014), and biophysical models (Honey et al. 2007; Breakspear 2017). By modeling brain networks as graphs, researchers employ graph-theoretic approaches to reveal key topological properties that underlie efficient communication and cognitive processes (Kan et al. 2022; Said et al. 2023). Understanding how these structural features relate to functional dynamics is a central goal in connectomics.
@@
To investigate the principles governing connectome formation and organization, generative models have been developed. Early approaches typically relied on a small set of predefined wiring rules or biological principles, such as cost-efficiency or growth mechanisms, to replicate observed network features in silico (Betzel et al. 2016; Kaiser and Hilgetag 2004; Henriksen, Pang, and Wronkiewicz 2016). While successful in capturing certain global properties, these rule-based methods often lack flexibility and depend on manually specified or constrained generative factors, limiting their ability to explore the full complexity of biological variability.
@@
More recently, deep generative models, such as Variational Autoencoders (VAEs), have emerged as powerful tools for analyzing and synthesizing complex data like brain networks (Yu et al. 2022; Zuo et al. 2023). These models are particularly well-suited for learning a low-dimensional latent representation of the network data, effectively performing dimensionality reduction. Furthermore, by manipulating or sampling from this learned latent space, these models enable the controlled synthesis of novel, biologically plausible network configurations, offering new avenues for exploring connectome variability and its functional implications (Tan et al. 2022; Dong et al. 2023). Prior works have also applied graph VAEs to macro-scale functional connectomes for discriminative tasks or developmental modeling (Behrouzi and Hatzinakos 2022).
Method
MICrONS Dataset and Preprocess
@@
We adopt the IARPA MICrONS dataset (Consortium et al. 2023), which encompasses a 1.4 × 0.87 × 0.84 mm volume of cortex from a P87 mouse, containing neurons within the area and their interconnections, which is the sole source for such high-resolution connectivity. Given the cortex’s division into six distinct layers, each associated with a specific level of information processing (Jones 2000; Felleman and Van Essen 1991; Reid and Alonso 1996), our study concentrates on cortical microcircuits organized as vertically oriented functional columns that traverse all laminar layers (Callaway 1998; Douglas and Martin 2004b). To model these microcircuits, we extracted cylindrical subvolumes oriented perpendicularly to the cortical layers. Each cylindrical unit comprises 80–100 neurons (both excitatory and inhibitory neurons), preserving intra-column connectivity while excluding connections extending beyond the columnar boundary. Our VAE was trained on 3,285 circuits from MICRONS dataset. The test set was from a spatially disjoint region to prevent data leakage.
Focusing only on network topology, we treat each circuit as a binary directed graph. We establish a canonical ordering rule, $\pi$, sorting neurons by their y-coordinates (reflecting cortical depth). This depth-based ordering is a deliberate choice, as it is both biologically significant (the y-coordinate inherently encodes the cortical layer) and technically crucial. Without this consistent ordering, determining the correct correspondence for element-wise comparison between original and generated adjacency matrices becomes intractable. Neuron type and synapse weight information are temporarily disregarded, as this work focuses solely on the topological structure.
We applied a consistent methodology to the fruit fly connectome. While the overall paradigm of using a generative model to learn a low-dimensional topological representation remains the same, the specific VAE architecture was adapted to accommodate the different structural properties of the fly brain.
Connectome Graph Variational Autoencoder
@@
As the goal is to find latent representations similar to “genomic bottleneck”, we need a generative model with “information bottleneck”. Thus, adopting a variational autoencoder (VAE) (Kingma and Welling 2022) on the connectome data can be formulated as follows: Assume a set of training connectome graphs $\mathbf{G}=\{\mathcal{G}_{\pi}^{(i)}\}$ is generated from the distribution of a set of unobserved latent representation $\textbf{z}=(z_{1},...,z_{m})$, where $z^{(i)}\sim p_{\theta^{*}}(z)$ and training data is sampled from true conditional $p_{\theta^{*}}(\mathcal{G}_{\pi}|z^{(i)})$. The data generation process is denoted by $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$. Following the standard VAE setting (Kingma and Welling 2022), we approximate the intractable posterior by $q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi})\approx p_{\theta}(\mathbf{z}|\mathcal{G}_{\pi})$ and maximize the evidence lower bound on the marginal likelihood of graph $\mathcal{G}^{(i)}$:
$\quad\log p_{\theta}(\mathcal{G}_{\pi}^{(i)})\geq\mathcal{L}(\phi,\theta;\mathcal{G}_{\pi}^{(i)})$
@@
$=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi}^{(i)})}\left[\log p_{\theta}(\mathcal{G}_{\pi}^{(i)}|\mathbf{z})\right]-\beta\mathrm{KL}\left[q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi}^{(i)})||p_{\theta}(\mathbf{z})\right],$
@@
where $\beta$ is a hyperparameter to balance the reconstruction loss and KL-divergence loss during the training process (Higgins et al. 2016).
@@
The proposed VAE model for connectome graphs comprises four main components. First, a node feature encoder employs a multi-layer, multi-head Graph Attention Network (GAT) (Veličković et al. 2018; Vaswani et al. 2023) to compute an embedding vector for each node. Second, a graph global encoder, which is a transformer encoder augmented with a special token (Devlin et al. 2019; Winter, Noé, and Clevert 2021; Vaswani et al. 2023), extracts an embedding for the entire sequence of node embeddings. This extracted sequence embedding serves as the graph-level embedding, upon which a 32-dimensional mean and variance are computed. The reparameterization of a standard VAE is applied to this 32D latent embedding representing the whole graph. Third, a node feature decoder, which is a transformer decoder, takes the graph-level embedding as input to reconstruct node features necessary for subsequent edge prediction. Finally, the edge predictor utilizes these decoded node features to predict the edges of the graph.
@@
The overall architecture of the VAE model is illustrated in Figure 1. The details of the 4 components are in Appendix B. The detailed encoder structure is depicted in Appendix Figure 8, and the detailed decoder structure is depicted in Appendix Figure 9.
@@
Controllable Connectome Generation by Sampling from the Latent Space
@@
When $\lambda\to\infty$, the energy model degenerates into a strictly constrained version (by $\mathcal{T}$). In this case we have:
@@
where the new normalization factor becomes $p(\Omega_{\mathcal{T}})=\int_{\Omega_{\mathcal{T}}}p(\mathbf{z})^{{1}/{\tau}}d\mathbf{z}$.
Reservoir Network with Structured Connectivity
We define reservoir network (Jaeger 2001) with a $D_{r}$ dimensional reservoir $\mathbf{R}$, input projection parameterized by $W_{in}\in\mathbb{R}^{D_{r}\times D}$ and output projection parameterized by $W_{out}\in\mathbb{R}^{D\times D_{r}}$. Reservoir $\mathbf{R}$ stores a reservoir state $\mathbf{r}(t)$ with dimension $D_{r}$, and the recurrent connection is represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{D_{r}\times D_{r}}$, whose elements are 0, 1 or -1, corresponding to non-connection, excitatory connection and inhibitory connection. We scale $\mathbf{A}$ to satisfy the pre-specified spectral radius $sr$. Given input vector $\mathbf{u}(t)$, the reservoir state is updated by $\mathbf{r}(t+1)=(1-\alpha)\mathbf{r}(t)+\alpha\cdot\tanh(\mathbf{A}\cdot\mathbf{r}(t)+W_{in}\cdot\mathbf{u}(t))$, where $\alpha$ is the leaking rate. $\mathbf{r}(t+1)$ is then flowed to the output coupler, where a mapping is done to obtain $\mathbf{v}(t+1)=W_{out}\cdot\mathbf{r}(t+1)$. During the training process, we kept the recurrent connection weights of the reservoir fixed, training only the input and readout layers. This approach allows us to study the computational functions that arise from the connection structure itself.
Experiments
Evaluation Metrics
We selected several graph theory metrics (descriptors) commonly employed in the analysis of connectomes (Rubinov and Sporns 2010), including mean degree, efficiency, transitivity, clustering coefficient, modularity, and assortativity. The detailed definitions of these metrics are provided in the Appendix C.
Microcircuit Reconstruction and Generation Results
Examples of original and reconstructed graphs obtained through the VAE are provided in the Appendix D. To validate the model’s generative capabilities, we evaluated several graph metrics on the generated samples. We employed the maximum mean discrepancy (MMD) (Gretton et al. 2012) to compare the distributions of these graph statistics between an equal number of generated and test graphs. Following the standard protocol established by GraphRNN (You et al. 2018), we specifically measured the distributions of degree, clustering coefficient and spectrum. For MMD computation, we utilized both the Gaussian Earth Mover’s Distance (EMD) kernel and the total variation (TV) distance (Liao et al. 2020). Given the absence of existing generative models specifically designed for connectome graphs, we selected three alternative models commonly used for molecule generation or general graph generation (Jo, Lee, and Hwang 2022; Xu et al. 2024; Chen et al. 2023) as baselines. The following Table 1 reports the MMD values for different graph metrics evaluated across these models and our proposed approach. To further check the diversity and novelty of the generated results, we verified that 2,000 generated samples are all unique among themselves and not present in the training data.
Figure 2 compares graphs generated by our proposed model and other models with the original data (column Ori.).
@@
GDSS (Jo, Lee, and Hwang 2022) 1.138 0.493 1.381 0.959 0.325 0.515
DisCo (Xu et al. 2024) 1.047 0.304 1.315 0.550 0.094 0.334
EDGE (Chen et al. 2023) 0.660 0.041 0.343 1.016 0.972 0.111
Ours $\boldsymbol{0.164}$ $\boldsymbol{0.028}$ $\boldsymbol{0.154}$ $\boldsymbol{0.021}$ $\boldsymbol{0.047}$ $\boldsymbol{0.007}$
Table 1: Demonstrating Model Effectiveness: Comparison with Baseline Models
@@
Figure 2: Generated Results Demonstrate the Authenticity of Our Model
Latent Representation Analysis
SHAP Analysis